Suppose that Randy Morgan is offered a job at $1600 per month with a guaranteed increase of $50 every six months for 5 yr. what will his salary be at the end of that time?
Answer by mananth(16949)

Initial salary = 1500
raise = $50 every six months
number oof years = 5
number of raises = 2*5 =10
It is an A P
a=1500, d=50, n= 10
ta = a+(n-1)d
=1500+(10-1)50
=1500+450
=1950
Income after 5 years = $1950